OpenStudy (anonymous):

Use either elimination or substitution to find the solutions (if any). 2y + 3x = 7 and 8y + 12x = 28

OpenStudy (anonymous):

Let's do this together! Ready?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh hi uber

OpenStudy (anonymous):

Equation 1 = (2y + 3x = 7) Equation 2 = (8y + 12x = 28) What I like to do first is mix the equations around to get them in the form of x= blahblah or y = blahblah. Equation1y = (2y + 3x = 7) 2y = -3x +7 y = (-3x + 7)/2 Equation1x = (2y + 3x = 7) 3x = 7-2y x = (7-2y)/3 Let me know when you're ready and I'll tell you what we should do next.

OpenStudy (anonymous):

Hi there, boog1!

OpenStudy (anonymous):

Just lemme know whenever you're ready to do the next step. Also, after we solve this, I'll also explain the elimination method.

OpenStudy (anonymous):

so which are we using elimination or substitution

OpenStudy (anonymous):

ok so now we take x=97-2y)/3 an and it to the second equation?

OpenStudy (anonymous):

That's right. For now we're doing substitution.

OpenStudy (anonymous):

9=(

OpenStudy (anonymous):

i kinda know that one

OpenStudy (anonymous):

8y + 12[(7-2y)/3]= 28 Looking good so far?

OpenStudy (anonymous):

yup i got that far

OpenStudy (anonymous):

Okie dokie. Now, this part is really ugly: 12[(7-2y)/3]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

What I would do is distribute the /3.

OpenStudy (anonymous):

So, [(7-2y)/3] becomes: 7/3 - 2y/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8y + 12[7/3 - 2y/3]= 28

OpenStudy (anonymous):

Now, we distribute the 12.

OpenStudy (anonymous):

8y + 84/3 - 24y/3 = 28 So far so good?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you see anything we can do to make the equation less complicated?

OpenStudy (anonymous):

8y+28-8y=28?

OpenStudy (anonymous):

Excellent.

OpenStudy (anonymous):

Hmm... an interesting conundrum...

OpenStudy (anonymous):

then we take the 8y out because there is 2?

OpenStudy (anonymous):

You see what our problem is, right?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Hahah.... I just noticed something funny.

OpenStudy (anonymous):

Equation 1 = (2y + 3x = 7) Equation 2 = (8y + 12x = 28) What I like to do first is mix the equations around to get them in the form of x= blahblah or y = blahblah. Equation1y = (2y + 3x = 7) 2y = -3x +7 y = (-3x + 7)/2 Equation1x = (2y + 3x = 7) 3x = 7-2y x = (7-2y)/3 Let me know when you're ready and I'll tell you what we should do next.

OpenStudy (anonymous):

So we end up with 0=0, right?

OpenStudy (anonymous):

so y=28

OpenStudy (anonymous):

NooooooooOOoooo :P

OpenStudy (anonymous):

"8y+28-8y=28?" 28 = 28 0=0

OpenStudy (anonymous):

You know what this means?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Hint: because we have no "X = something" or "Y = something", this means one of two choices: "Unlimited answers" or... "Not a real equation"

OpenStudy (anonymous):

ok this is the next question for this problem . For equation ___________ 1. The x-intercept is ___. 2. The y-intercept is ___. 3. The slope is ___.

OpenStudy (anonymous):

The way to find this out is to see whether the last sentence we're left with is true or not. So... if you ever got something like 0 = 5, you know that's a false, so there are 'no solutions" So tell me what our answer is.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

this is unlimited answers

OpenStudy (anonymous):

It's not zero, though. Remember, it's only 0 if it says "y=0" or "x=0". What we got was "8y-8y+28=28". Which turns into "28=28", which you can turn into "0=0"

OpenStudy (anonymous):

Excellent! Unlimited answers. Now, if it's unlimited answers, what does this tell us about the two lines?

OpenStudy (anonymous):

they will be paraell

OpenStudy (anonymous):

If they're parallel, this actually means "no answer exists". An "answer" is when the two lines touch. Parallel lines never touch.

OpenStudy (anonymous):

i am looking at my notes

OpenStudy (anonymous):

Think of it this way: Unlimited answers means the lines are always on top of each other, right? And in a two dimensional world, what happens if two lines are on top of each other?

OpenStudy (anonymous):

unique solution

OpenStudy (anonymous):

Unique solution would mean that they hit only once. Let's do a quick twiddla thingy.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://www.twiddla.com/584279